x-3y=-8
y+2=1
x+42=7
How do I solve this problem?

Answer by Alan3354(69443)

x-3y=-8
y+2=1
x+42=7
How do I solve this problem?
------------
y+2=1
y = -1
---------
x+42=7
x = -35
-------
The 1st eqn doesn't fit.